# Predicting Forced Responses of Probability Distributions via the Fluctuation–Dissipation Theorem and Generative Modeling

### v1

@@ Abstract
We present a novel data-driven framework for estimating the response of higher-order moments of nonlinear stochastic systems to small external perturbations. The classical Generalized Fluctuation–Dissipation Theorem (GFDT) links the unperturbed steady-state distribution to the system’s linear response. Standard implementations rely on Gaussian approximations, which can often accurately predict the mean response but usually introduce significant biases in higher-order moments—such as variance, skewness, and kurtosis. To address this limitation, we combine GFDT with recent advances in score-based generative modeling, which enable direct estimation of the score function from data without requiring full density reconstruction. Our method is validated on three reduced-order stochastic models relevant to climate dynamics: a scalar stochastic model for low-frequency climate variability, a slow–fast triad model mimicking key features of the El Niño-Southern Oscillation (ENSO), and a six-dimensional stochastic barotropic model capturing atmospheric regime transitions. In all cases, the approach captures strongly nonlinear and non-Gaussian features of the system’s response, outperforming traditional Gaussian approximations.
@@
Predicting how complex stochastic systems respond to small external perturbations is central in physics, climate science, and engineering. We combine the Generalized Fluctuation–Dissipation Theorem with score‑based generative modeling (KGMM) to accurately capture mean and higher‑order (variance, skewness, kurtosis) responses—even in strongly non‑Gaussian regimes.
@@ 1 Introduction
Understanding how a physical system responds to external perturbations is a central challenge in physics, climate science, and engineering (Strogatz, 2018; Palmer, 2001). In many applications, one is not only interested in the shift of the system’s mean state under a small forcing but also in changes in its higher-order moments—such as variance, skewness, and kurtosis. Capturing responses in higher-order moments is crucial for reconstructing the perturbed steady-state probability distribution and for building a principled, causal framework to analyze how perturbations shape the system’s overall response, including distributional tails (Majda et al., 2010b). For instance, in climate science, assessing the full redistribution of probability (e.g., changes in the frequency of heat waves or cold spells) is as crucial as estimating an average temperature increase (IPCC, 2013).
@@
A key theoretical tool in this context is the Generalized Fluctuation-Dissipation Theorem (GFDT), which relates the steady-state probability density function (PDF) of a system to its linear response under a time-dependent external perturbation (Giorgini et al., 2024b; Cooper and Haynes, 2011; Baldovin et al., 2020; Ghil and Lucarini, 2020; Majda et al., 2008b, 2005). In principle, GFDT offers a route to predict how the entire distribution of an observable changes without explicitly perturbing the system. However, practical application of the GFDT is hindered by the difficulty of accurately estimating the full, often high-dimensional and non-Gaussian, steady-state PDF. In many cases the Gaussian approximation—where the system’s PDF is replaced by a Gaussian distribution sharing the same mean and covariance—is employed. It has been shown empirically that the Gaussian approximation has high-skill in predicting responses in the mean even in nonlinear systems (e.g. Gritsun and Branstator (2007); Gershgorin and Majda (2010); Baldovin et al. (2020)) but introduces systematic biases for higher-order moments (Majda et al., 2008c; Majda and Qi, 2019).
@@
Recent advances in data-driven methodologies offer promising alternatives. In particular, score-based generative modeling has emerged as a powerful approach for sampling from complex distributions by learning the gradient of the log-PDF (the score function) directly from data, thereby circumventing the need for full density estimation (Song et al., 2021). Moreover, modern clustering-based algorithms have shown that statistical properties of high-dimensional systems can be estimated more reliably from coarse-grained observational data than the systems’ detailed dynamical trajectories (Falasca et al., 2024, 2025; Giorgini et al., 2025b, 2024a; Souza, 2024a, b; Souza and Silvestri, 2024).
@@
These data-driven advances are particularly valuable in settings where directly solving the full partial differential equations (PDEs) governing a system’s dynamics is computationally prohibitive or even unknown. This limitation has motivated the development of reduced-order models, which aim to reproduce the essential statistical behavior of complex systems at a fraction of the computational cost. In climate science, in particular, stochastic reduced-order models have played a central role in both understanding and forecasting system behavior (Hasselmann, 1976; Penland, 1989; Penland and Sardeshmukh, 1995; Majda et al., 1999, 2001; Kondrashov et al., 2015; Strounine et al., 2010; Lucarini and Chekroun, 2023). For example, reduced Markovian models have been successfully employed to derive effective dynamics for slow variables in multiscale systems (Kravtsov et al., 2005; Majda et al., 2008a; Chen et al., 2019; Keyes et al., 2023; Giorgini et al., 2022). These approaches facilitate our understanding of high-dimensional dynamical systems by focusing on core processes while also improving computational efficiency. Furthermore, reduced-order models also shed light on how non-Gaussian features and intermittent behavior emerge from interactions between resolved and unresolved scales.
@@
Building on these ideas, recent studies have combined the GFDT framework with generative score-based techniques to improve predictions of the mean response in high-dimensional systems (Giorgini et al., 2024b). In this paper, we build on the framework presented in (Giorgini et al., 2024b) by constructing higher‑order response functions for nonlinear reduced‑order stochastic models relevant to climate science. Specifically, we examine the response of the mean (first moment) and the second, third, and fourth central moments of an observable to small external perturbations. By incorporating information from these higher-order moments, our approach provides an accurate estimate of the perturbed steady-state PDF, which is crucial for understanding changes in variability and the occurrence of extreme events. We test the approach on the following reduced-order stochastic models of increasing complexity: a one-dimensional stochastic model for low‑frequency climate variability, a slow–fast triad model designed to capture key features of the El Niño-Southern Oscillation (ENSO), and a six-dimensional stochastic barotropic model that reproduces atmospheric regime transitions (Charney and DeVore, 1979; De Swart, 1988; Crommelin et al., 2004). Our results demonstrate that the new method can accurately reproduce the steady-state PDFs and outperforms traditional Gaussian approximations in capturing higher-order moment responses.
@@
The remainder of the paper is organized as follows. In Section 2, we review the derivation and practical implementation of the GFDT and we detail the extraction of the score function from data through a recently proposed clustering-based algorithms. In Section 3, we present numerical experiments on the three reduced-order models, highlighting the improved skill of our approach in capturing higher-order responses compared to the standard Gaussian approximation. Section 4 focuses on limitations and practical considerations of our proposal in the presence of partial observations from high-dimensional dynamical systems. Finally, Section 5 summarizes our findings and discusses directions for future work.
@@ 2 Methods
In this section, we present the theoretical and computational foundations of our approach. We begin by deriving the Generalized Fluctuation-Dissipation Theorem (GFDT), which forms the basis for computing response functions in nonlinear stochastic systems. We describe the KGMM (K‐means Gaussian Mixture Modeling) algorithm from (Giorgini et al., 2025a) for data-driven estimation of the score function, which plays a central role in implementing the GFDT for this work. Additional details on the maximum entropy framework used to reconstruct perturbed steady-state distributions from estimated moment responses are provided in the Appendix.
@@ 2.1 The Generalized Fluctuation-Dissipation Theorem
where $\bm{x}\in\mathbb{R}^{n}$, $\bm{F}(\bm{x})$ is a deterministic drift term, $\bm{\sigma}$ is the noise amplitude matrix, and $\bm{\xi}$ is a standard Gaussian white noise process. The system evolves towards a steady state described by a probability density $\rho_{S}(\bm{x})$, which satisfies the stationary Fokker–Planck equation:
@@
with $\bm{D}=\bm{\sigma}\bm{\sigma}^{\top}$ the diffusion matrix, and :::\,: denoting the double contraction of tensors.
@@
By formally integrating this linearized equation, the resulting first-order correction to the expectation value of any observable $A(\bm{x})$ is given by
@@
Note that climate studies traditionally focus on the specific case of changes in the external forcing as $\bm{u}(\bm{x})_{i}=\bm{e}_{i}$, further simplifying Eq. 10 to $B(\bm{x})=-\nabla\ln\rho_{S}(\bm{x})$.
@@
which yields
@@
In this case, it is straightforward to verify that, under external forcings $\bm{u}(\bm{x})_{i}=\bm{e}_{i}$ the response function reduces to $\bm{R}(t)=\langle A(\bm{x}(t))\bm{\Sigma}^{-1}(\bm{x}-\bm{\mu})(0)\rangle_{0}$, as commonly adopted in climate studies. Note that while the analytical score is approximated using that of a Gaussian distribution, the ensemble averages $\langle\cdot\rangle_{0}$ are still obtained using the original data and therefore over the system’s steady state distribution. This approach differs from linear inverse model strategies (e.g., Penland (1989)), which assume a time-invariant Gaussian measure from the outset. Such models would yield zero response to any quadratic functional, such as the variance (see Majda et al. (2010a)). Because of this distinction, the approximation discussed here is sometimes referred to as a quasi-Gaussian approximation in the literature. For simplicity, we will simply refer to it as the Gaussian approximation throughout this work.
@@ 2.2 Score Function Estimation via KGMM
$p(\bm{x})=\frac{1}{N}\sum_{i=1}^{N}\mathcal{N}(\bm{x}\mid\bm{\mu}_{i},\sigma^{2}\bm{I}),$ (15)
where the $\bm{\mu}_{i}$ are data samples drawn from the steady-state distribution $\rho_{S}(\bm{x})$, and $\sigma^{2}$ is the (isotropic) covariance amplitude of the Gaussian kernels. The corresponding score function reads
$\nabla\ln p(\bm{x})=-\frac{1}{\sigma^{2}}\sum_{i=1}^{N}\frac{\mathcal{N}(\bm{x}\mid\bm{\mu}_{i},\sigma^{2}\bm{I})(\bm{x}-\bm{\mu}_{i})}{p(\bm{x})}.$ (16)
While Eq. (16) provides a direct expression for the score, it becomes numerically unstable for small $\sigma$, as the density and its derivative become highly sensitive to local data fluctuations.
KGMM circumvents this issue by exploiting a probabilistic identity: define $\bm{x}=\bm{\mu}+\sigma\bm{z}$ with $\bm{z}\sim\mathcal{N}(0,\bm{I})$. Giorgini et al. (2025a) derived the following relation in the continuous-data limit:
$\nabla\ln p(\bm{x})=-\frac{1}{\sigma^{2}}\mathbb{E}[\bm{z}\mid\bm{x}],$ (17)
i.e., the score function is the conditional expectation of the kernel displacements $\bm{z}$, rescaled by $\sigma^{2}$.
This identity is computed empirically using the following steps:
  1. 1.
    Generate perturbed samples $\bm{x}_{i}=\bm{\mu}_{i}+\sigma\bm{z}_{i}$, where $\bm{\mu}_{i}$ are the original data points and $\bm{z}_{i}\sim\mathcal{N}(0,\bm{I})$.
  2. 2.
    Partition the perturbed samples $\{\bm{x}_{i}\}$ into $N_{C}$ clusters $\{\Omega_{j}\}$ using bisecting K-means clustering.
  3. 3.
  4. 4.
    Estimate the score function at the centroid $\bm{C}_{j}$ as
    $\nabla\ln\rho_{S}(\bm{C}_{j})\approx-\frac{1}{\sigma^{2}}\mathbb{E}[\bm{z}\mid\bm{x}\in\Omega_{j}].$ (19)
  5. 5.
The number of clusters $N_{C}$ must therefore be chosen carefully to balance the trade-off between resolution and noise. Empirically, a useful scaling relation is
$N_{C}\propto\sigma^{-d},$ (20)
where $d$ is the effective dimensionality of the dataset and $\sigma$ is the kernel width. This scaling ensures that clusters remain small enough to capture local gradient structure while still containing enough points for robust averaging.
The choice of $\sigma$ plays a central role in the KGMM algorithm. Small values of $\sigma$ yield estimates of the score function that are closer to the one associated with the true steady-state distribution, as the perturbation introduced by the convolution kernel becomes negligible. However, this comes at the cost of increased statistical noise, since the displacements become more sensitive to sample variability. Conversely, larger values of $\sigma$ smooth out the fluctuations, leading to more stable estimates but of a score function associated with a more strongly perturbed distribution. The optimal value of $\sigma$ thus balances these competing effects—reducing bias while maintaining statistical reliability.
To interpolate the discrete score estimates $(\bm{C}_{j},\nabla\ln\rho_{S}(\bm{C}_{j}))$, we train a fully connected feedforward neural network with two hidden layers, using the Swish activation function between layers and a linear activation on the output layer. The models and training parameters are:
All networks are trained using the Adam optimizer with mean‐squared‐error loss on the predicted score.

### v2

@@ Abstract
We present a novel and flexible data-driven framework for estimating the response of higher-order moments of nonlinear stochastic systems to small external perturbations. The classical Generalized Fluctuation–Dissipation Theorem (GFDT) links the unperturbed steady-state distribution to the system’s linear response. While standard implementations relying on Gaussian approximations can predict the mean response, they often fail to capture changes in higher-order moments. To overcome this, we combine GFDT with score-based generative modeling to estimate the system’s score function directly from data. We demonstrate the framework’s versatility by employing two complementary score estimation techniques tailored to the system’s characteristics: (i) a clustering-based algorithm (KGMM) for systems with low-dimensional effective dynamics, and (ii) a denoising score matching method implemented with a U-Net architecture for high-dimensional, spatially-extended systems where reduced-order modeling is not feasible. Our method is validated on several stochastic models relevant to climate dynamics: three reduced-order models of increasing complexity and a 2D Navier-Stokes model representing a turbulent flow with a localized perturbation. In all cases, the approach accurately captures strongly nonlinear and non-Gaussian features of the system’s response, significantly outperforming traditional Gaussian approximations.
@@
Predicting how complex stochastic systems respond to small external perturbations is central in physics, climate science, and engineering. We combine the Generalized Fluctuation–Dissipation Theorem with score-based generative modeling to accurately capture mean and higher-order (variance, skewness, kurtosis) responses. Our framework demonstrates significant flexibility, providing accurate results for both well-behaved reduced-order models and for high-dimensional turbulent systems where such reductions are not possible. This work provides a powerful, data-driven tool for understanding how small forcings alter the full probability distribution, including the statistics of extreme events, in strongly non-Gaussian systems.
@@ 1 Introduction
Understanding how a physical system responds to external perturbations is a central challenge in physics, climate science, and engineering (Strogatz, 2018; Palmer, 2001; Baiesi and Maes, 2013). In many applications, one is not only interested in the shift of the system’s mean state under a small forcing but also in changes in its higher-order moments—such as variance, skewness, and kurtosis. Capturing responses in higher-order moments is crucial for reconstructing the perturbed steady-state probability distribution and for building a principled, causal framework to analyze how perturbations shape the system’s overall response, including distributional tails (Majda et al., 2010b). For instance, in climate science, assessing the full redistribution of probability (e.g., changes in the frequency of heat waves or cold spells) is as crucial as estimating an average temperature increase (IPCC, 2013).
@@
A key theoretical tool in this context is the Generalized Fluctuation-Dissipation Theorem (GFDT), which relates the steady-state probability density function (PDF) of a system to its linear response under a time-dependent external perturbation (Giorgini et al., 2024b; Cooper and Haynes, 2011; Baldovin et al., 2020; Ghil and Lucarini, 2020; Majda et al., 2008b, 2005). In principle, GFDT offers a route to predict how the entire distribution of an observable changes without explicitly perturbing the system. However, practical applications of the GFDT are hindered by the difficulty of accurately estimating the full, often high-dimensional and non-Gaussian, steady-state PDF. In many cases the Gaussian approximation—where the system’s PDF is replaced by a Gaussian distribution sharing the same mean and covariance—is employed. It has been shown empirically that the Gaussian approximation has high-skill in predicting responses in the mean even in nonlinear systems (e.g. Gritsun and Branstator (2007); Gershgorin and Majda (2010); Baldovin et al. (2020)) but introduces systematic biases for higher-order moments (Majda et al., 2008c; Majda and Qi, 2019).
@@
Recent advances in data-driven methodologies offer promising alternatives. For instance, operator-theoretic frameworks connect a system’s response to the spectral properties of its evolution operator. This viewpoint allows for expressing the response formula in terms of the poles and residues of the susceptibility function of the unperturbed system (Gutiérrez and Lucarini, 2022), and it has led to practical data-driven methods for constructing response operators from time series (Chekroun et al., 2024). More generally, response theory for any Markovian system can be formulated via a fluctuation-dissipation relation for the process generator (Lucarini, 2025). In particular, score-based generative modeling has emerged as a powerful approach for sampling from complex distributions by learning the gradient of the log-PDF (the score function) directly from data, thereby circumventing the need for full density estimation (Song et al., 2021). Moreover, modern clustering-based algorithms have shown that statistical properties of high-dimensional systems can be estimated more reliably from coarse-grained observational data than the systems’ detailed dynamical trajectories (Falasca et al., 2024, 2025; Giorgini et al., 2025b, 2024a; Souza, 2024a, b; Souza and Silvestri, 2024).
@@
These data-driven advances are particularly valuable in settings where directly solving the full partial differential equations (PDEs) governing a system’s dynamics is computationally prohibitive or even unknown. This limitation has motivated the development of reduced-order models, which aim to reproduce the essential statistical behavior of complex systems at a fraction of the computational cost. In many high-dimensional dynamical systems, the vast majority of the system’s energy and variability is often concentrated in a small number of large-scale, coherent patterns, commonly identified through empirical orthogonal functions (EOFs) or other dimensionality reduction techniques. These dominant patterns capture the essential dynamics of the system while filtering out higher-frequency, smaller-scale fluctuations. One can identify these dominant patterns and build a stochastic model for their dynamics alone, effectively reducing the dimensionality of the problem while retaining the key physical processes that govern the system’s behavior. In climate science, in particular, stochastic reduced-order models have played a central role in both understanding and forecasting system behavior (Hasselmann, 1976; Penland, 1989; Penland and Sardeshmukh, 1995; Majda et al., 1999, 2001; Kondrashov et al., 2015; Strounine et al., 2010; Lucarini and Chekroun, 2023). For example, reduced Markovian models have been successfully employed to derive effective dynamics for slow variables in multiscale systems (Kravtsov et al., 2005; Majda et al., 2008a; Chen et al., 2019; Keyes et al., 2023; Giorgini et al., 2022; Giorgini, 2025).
Similar reduced-order modeling approaches have proven essential across diverse research areas dealing with nonequilibrium systems, including turbulent fluid dynamics (Berkooz et al., 1993; Rowley and Dawson, 2017), plasma physics (Peerenboom et al., 2015; Farcaş et al., 2024), biological systems (Brunton et al., 2016b; Champion et al., 2019), chemical reaction networks (Valorani et al., 2006; Lu and Law, 2008), and neuroscience (Brunton et al., 2016a; Proctor et al., 2016).
These approaches facilitate our understanding of high-dimensional dynamical systems by focusing on core processes while also improving computational efficiency. Furthermore, reduced-order models also shed light on how non-Gaussian features and intermittent behavior emerge from interactions between resolved and unresolved scales, (Falasca, 2025).
However, constructing accurate response functions for reduced-order models (ROMs) presents inherent limitations. For the GFDT to yield an accurate response from a ROM, a specific set of conditions should be met: (i) the retained slow modes must be dynamically well separated from the fast, truncated modes which will be treated as noise; and (ii) the external perturbation should project primarily onto the slow, coherent modes included in the ROM. When these conditions are violated, the ROM may fail to capture the full system’s response to perturbations. Nevertheless, many problems of practical interest fall into this favorable scenario, such as studying the climate’s response to large-scale forcings (e.g., uniform CO2 increase) that naturally align with the system’s dominant coherent modes (Marvel et al., 2016).
@@
Building on these ideas, recent studies have combined the GFDT framework with generative score-based techniques to improve predictions of the mean response in high-dimensional systems (Giorgini et al., 2024b). In this paper, we build on the framework presented in (Giorgini et al., 2024b) by constructing higher‑order response functions for nonlinear reduced‑order stochastic models relevant to climate science. Specifically, we examine the response of the mean (first moment) and the second, third, and fourth central moments of an observable to small external perturbations. By incorporating information from these higher-order moments, our approach provides an accurate estimate of the perturbed steady-state PDF, which is crucial for understanding changes in variability and the occurrence of extreme events. We test the approach on the following reduced-order stochastic models of increasing complexity: a one-dimensional stochastic model for low‑frequency climate variability, a slow–fast triad model designed to capture key features of the El Niño-Southern Oscillation (ENSO), and a six-dimensional stochastic barotropic model that reproduces atmospheric regime transitions (Charney and DeVore, 1979; De Swart, 1988; Crommelin et al., 2004). We also study the case where the applied perturbation couples with all modes of the high-dimensional system. In this latter case, approximating the system’s behaviour with a ROM is not possible. We show how our method can be generalized to this case by considering as an example a 2D Navier-Stokes model with a localized perturbation. Our results demonstrate that the new method can accurately reproduce the steady-state PDFs and outperforms traditional Gaussian approximations in capturing higher-order moment responses.
@@
The remainder of the paper is organized as follows. In Section 2, we review the derivation and practical implementation of the GFDT and introduce the two complementary, data-driven techniques used for score function estimation: a clustering-based algorithm (KGMM) for low-dimensional systems, and a denoising score-matching approach for high-dimensional systems. To keep the main text focused, the methodological details are outlined in the Appendix. In Section 3, we present numerical experiments on several models of increasing complexity—the three reduced-order climate models and a high-dimensional 2D Navier-Stokes model—highlighting the improved skill of our approach in capturing higher-order responses compared to the standard Gaussian approximation. Section 4 discusses the practical considerations and limitations of the framework, contrasting the conditions under which reduced-order models are applicable with scenarios requiring direct estimation on the full state space. Finally, Section 5 summarizes our findings and discusses directions for future work.
@@ 2 Methods
In this section, we present the theoretical and computational foundations of our approach. We begin by deriving the Generalized Fluctuation-Dissipation Theorem (GFDT), which forms the basis for computing response functions in nonlinear stochastic systems. We then present two complementary, data-driven techniques for estimating the score function, which is central to implementing GFDT: (i) the KGMM (K-means Gaussian Mixture Modeling) algorithm (Giorgini et al., 2025a), used for low-dimensional systems (our scalar, triad, and barotropic ROMs); and (ii) denoising score matching, used for higher-dimensional systems (e.g., the 2D Navier–Stokes vorticity field on a $32\times 32$ grid). Throughout this work we follow the rule of thumb: lower-dimensional in state space $\rightarrow$ KGMM with a multilayer perceptron architecture; higher-dimensional in state space $\rightarrow$ denoising score-matching with a U -Net architecture. See Appendix § 6.3 for detailed guidance and a practical comparison. Additional details on the maximum entropy framework used to reconstruct perturbed steady-state distributions from estimated moment responses are provided in the Appendix (Appendix § 1).
@@ 2.1 The Generalized Fluctuation-Dissipation Theorem
where $\bm{x}\in\mathbb{R}^{n}$, $\bm{F}(\bm{x})$ is a deterministic drift term, $\bm{\sigma}$ is the noise amplitude matrix, and $\bm{\xi}$ is a standard Gaussian white noise process. The noise amplitude matrix $\bm{\sigma}$ can be constant, corresponding to additive noise, or a state-dependent function $\bm{\sigma}(\bm{x})$, corresponding to multiplicative noise. In the latter case, the equation is interpreted in the Itô sense, a choice reflected in the form of the Fokker-Planck operator below. Our framework relies on the assumption that the system is ergodic, admitting a unique and smooth invariant measure $\rho_{S}(\bm{x})$. Sufficient conditions for this are, for instance, the hypoellipticity of the Fokker-Planck operator and the existence of a Lyapunov function that ensures the confinement of trajectories (Pavliotis, 2014).
The system evolves towards a steady state described by a probability density $\rho_{S}(\bm{x})$, which satisfies the stationary Fokker–Planck equation:
@@
with $\bm{D}=\bm{\sigma}\bm{\sigma}^{\top}$ the diffusion matrix, and ::\, denoting the double contraction of tensors.
@@
For the perturbed density $\rho_{S}+\delta\rho$ to remain a valid probability distribution, the total probability must be conserved. This requires the integral of the density variation over the entire state space to be zero at all times. This consistency constraint, $\int\delta\rho(\mathbf{x},t)\,d\mathbf{x}=0$, is satisfied by Equation (5), given that the Fokker-Planck operator $\mathcal{L}_{0}$ preserves normalization (i.e., its integral over the domain is zero, assuming no-flux boundary conditions) and the system is unperturbed at $t=0$ (see Appendix § 2.)
By formally integrating this linearized equation (see Appendix § 3), the resulting first-order correction to the expectation value of any observable $A(\bm{x})$ is given by
@@
The structure of the conjugate observable $B(\bm{x})$ depends critically on the nature of the perturbation $\bm{u}(\bm{x})$. The expression simplifies considerably for a state-independent forcing, where $\bm{u}(\bm{x})$ is a constant vector. In this scenario, often employed in idealized models (Majda et al., 2010a) or certain diagnostic frameworks (Penland and Sardeshmukh, 1995), the divergence term vanishes ($\nabla\cdot\bm{u}(\bm{x})=0$), and $B(\bm{x})$ becomes proportional to the score function, $\nabla\ln\rho_{S}(\bm{x})$. In contrast, physically realistic forcings in climate science cab be inherently state-dependent. For example, the climatic forcing associated with an increase in CO2 concentration depends non-linearly on the state of the atmosphere (e.g., its temperature and water vapor content), meaning $\bm{u}=\bm{u}(\bm{x})$. For such cases, the divergence term $\nabla\cdot\bm{u}(\bm{x})$ is non-zero and physically crucial, making the full expression for $B(\bm{x})$ essential for an accurate response calculation (Lucarini et al., 2017; Lembo et al., 2020).
@@
where $\bm{\mu}$ and $\bm{\Sigma}$ respectively represent the mean and covariance matrix of the system. In this case we have:
@@
In this case, it is straightforward to verify that, under external forcings $\bm{u}(\bm{x})_{i}=\bm{e}_{i}$ the response function reduces to $\bm{R}(t)=\langle A(\bm{x}(t))\bm{\Sigma}^{-1}(\bm{x}-\bm{\mu})(0)\rangle_{0}$, as commonly adopted in climate studies Majda et al. (2005). Note that while the analytical score is approximated using that of a Gaussian distribution, the ensemble averages $\langle\cdot\rangle_{0}$ are still obtained using the original data and therefore over the system’s steady state distribution. This approach differs from linear inverse model strategies (e.g., Penland (1989)), which assume a time-invariant Gaussian measure from the outset. Such models would yield zero response to any quadratic functional, such as the variance (see Majda et al. (2010a); Falasca (2025)). Because of this distinction, the approximation discussed here is sometimes referred to as a quasi-Gaussian approximation in the literature. For simplicity, we will simply refer to it as the Gaussian approximation throughout this work.
@@ 2.2 Score Function Estimation via KGMM
To apply the Generalized Fluctuation-Dissipation Theorem (GFDT), we need the score function, $\nabla\ln\rho_{S}(\bm{x})$, which is the gradient of the logarithm of the system’s steady-state distribution. Since this function is rarely known analytically, it must be estimated from data. For systems where the dynamics can be effectively described in a low-dimensional space (as in our scalar, triad, and barotropic ROMs), we employ the KGMM (K-means Gaussian Mixture Modeling) algorithm, a hybrid statistical-learning method designed for efficient and accurate score estimation (Giorgini et al., 2025a).
The KGMM method computes the score by leveraging a probabilistic identity that avoids the numerical instabilities associated with direct differentiation of a data-based density estimate. The procedure begins by creating a new set of “perturbed” samples, where each original data point is shifted by a small, random vector drawn from a Gaussian distribution. This construction is closely connected to denoising score matching: training a network on perturbed samples to predict the added noise is equivalent to learning the score of a Gaussian–smoothed density (i.e., a Gaussian mixture with isotropic kernels); see (Kamb and Ganguli, 2024; Baptista et al., 2025). The algorithm’s core insight is that the score function is directly proportional to the conditional expectation—or local average—of these displacements. In practice, this is achieved by partitioning the perturbed data samples into clusters and computing the average displacement within each cluster, which provides a discrete estimate of the score at the cluster centroids. Finally, a neural network is trained to interpolate these discrete estimates, yielding a continuous representation of the score function over the entire domain.
Crucially, both KGMM and denoising score-matching is introduced to avoid the numerical instabilities of the direct GMM score estimator—differentiating a Gaussian-mixture density becomes ill-conditioned and noise-amplifying as the kernel width $\sigma_{G}\to 0$—by recasting score evaluation as a stable conditional-expectation regression (see Appendix § 6.1).
A detailed description of the algorithm, including the governing equations, implementation steps, and hyperparameter choices, is provided in Appendix § 6.1.
2.3 Score Function Estimation via Denoising Score-Matching
When a system’s dynamics cannot be effectively reduced to a low-dimensional model—for instance, if there is no clear separation of timescales or if a perturbation couples with all modes of a high-dimensional system—the KGMM clustering algorithm becomes computationally prohibitive. In these scenarios, we estimate the score function directly on the full, high-dimensional state space.
For this purpose, we employ a denoising score matching approach, training a neural network to learn the score directly from data (Karras et al., 2022). This method circumvents explicit clustering by framing the problem as one of learning to denoise perturbed samples. We use a U-Net architecture, which is particularly well-suited for spatially-extended systems due to its encoder-decoder structure with skip connections that preserve spatial information across multiple scales. The network is trained to minimize a denoising score matching loss function. We use the same methodology as what was implemented in (Giorgini et al., 2024b) with appropriate modifications to higher-order responses. This technique provides a scalable and powerful alternative for score estimation in high-dimensional settings where reduced-order modeling is not applicable. The full details of the U-Net architecture, loss function, and training procedure are available in Appendix § 6.2.